420lol:

Herman opened a credit card at a department store with an APR of 24.45% compounded semiannually. what is the APY on this credit card?

Vocaloid:

@sillybilly123 whenever you get a chance

ronndjaia:

Semiannually means two. so add 24.45 + 24.45 the answer would be A) 48.90%

sillybilly123:

i think these are American ideas, i'd be amazed if the EU allowed compounding to bend the meaning of APR: makes absolutely no sense. So, having googled this, and for semi-annual: \(APY = \left( 1 + \dfrac{APR}{2}\right)^2 - 1 = \left( 1 + \dfrac{0.2445}{2}\right)^2 - 1 = \dots\)